OpenStudy (anonymous):

write the equation of the line whose slope is 3 and passes through the point (1,4) put final answer in slope intercept form?

OpenStudy (anonymous):

y = m*x + b is the slope intercept form (It is formula of a straight line) (m is the slope (m=3), x=1 and y =4 are given values) Therefore b can be calculated as 4 = 3 * 1 + b b=1 So the slope intercept form is y = m * x + b y = 3 *x + 1 Hope this helps :)